Grade 12th passMagical Mathematics[Interesting Approach]

find the greatest six digit number that exactly divisible by 27 45 60 72 and 96

ANIL KUMAR
8 Years agoGrade 12th pass

1 Answer

Gulshan Goyal
8 Years ago
lcm of given numbers is 4320
greatest 6 digit no. can be calculated as  [1000000/4320]=231,
Answer=231*4320=997920